# Networks of neural networks: more is different

### v1

@@ 4 A performance-driven revision
Figure 9: We consider a dataset consisting of 6 digits, each represented by $58\times 52$ pixels, and analyze a three-layer network governed by the Hamiltonian (34) with parameters $N=3016$, $K=6$, $H=0.2$, and $\lambda=1.3$. A mixture of digits $0,1,6$ is then prepared and presented as input to each layer of the network, which is subsequently updated through MC simulations with $\beta=4$. The upper part of the figure illustrates the evolution of the neuronal configurations $\bm{\sigma}^{1}$, $\bm{\sigma}^{2}$, and $\bm{\sigma}^{3}$. Correspondingly, the lower part displays the evolution of the associated Mattis magnetizations $\bm{m}^{1}$, $\bm{m}^{2}$, and $\bm{m}^{3}$. Different colors are used to distinguish the magnetizations related to the different patterns comprising the dataset, with emphasis on the digits included in the input mixture, as highlighted in the legend.
@@ 5 Conclusions
In particular, here, given a dataset of binary vectors $\bm{\xi}=\{\bm{\xi}^{\mu}\}_{\mu=1,...,K}\in\{-1,+1\}^{N\times K}$, the input is given by mixtures like $\bm{\sigma}^{(h)}=\textrm{sign}(\bm{\xi}^{1}+\bm{\xi}^{2}+...+\bm{\xi}^{L})$ – and this is interpreted as the initial neuronal configuration for each layer, that is, $\bm{\sigma}^{a}=\bm{\sigma}^{(h)}$ for $a=1,...,L$ – while the desired output is given by $\bm{\sigma}^{(1,2,...,L)}:\bm{\sigma}^{\ell}=\bm{\xi}^{\ell}$ for $\ell=1,...,L$ (without loss of generality) – and this is interpreted as the target equilibrium state reached by the system.
@@
We emphasize that the kind of interactions implemented in this network yields a plethora of minima which can impair the disentanglement of the neuronal configuration $\bm{\sigma}^{(h)}$ into $\bm{\sigma}^{(1,2,...,L)}$. A way to see this is by considering an equivalent model obtained by applying a Hubbard-Stratonovich transformation to the model’s partition function (see App. A) and notice that the interaction among the dummy variables $z$’s is characterized by a high degree of frustration, especially compared with other layered associative-memory models, see e.g., AABCR-2024 . Many possible adjustments can be implemented to improve the performance of this model, for instance one can revise the Hebbian kernel to obtain a projection kernel Kohonen-1972 ; FAB-NN2019 that reduces the detrimental effects due to interference among the stored patterns, or allow for higher-order interactions FAB-NN2019 ; Krotov2018 ; Pedreschi1 ; Pedreschi2 which make the desired minima more stable. The last strategy is implemented in the last part of the paper, specifically, the pair-wise heterogeneous terms, namely those involving neurons from different layers, are replaced by fourth-order terms. Remarkably, this revision makes the model invariant under the spin-flip of a single layer, yields to more attractive disentangled states and therefore to a better performing model as far as the disentanglement task is concerned.
@@ Appendix A RS solution by interpolation technique
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{\xi},\bm{h})&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{(a)}\}}\exp\Bigg{[}\dfrac{\beta N}{2}\displaystyle\sum\limits_{\mu=1}^{K}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}\Bigg{]}.\end{array}$ (A.2)
@@
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{\xi},\bm{h})&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{(a)}\}}\exp\Bigg{[}\dfrac{\beta N}{2}\displaystyle\sum\limits_{\mu=1}^{L}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\dfrac{\beta N}{2}\displaystyle\sum\limits_{\mu>L}^{K}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}\Bigg{]}\end{array}$ (A.3)
@@
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{\xi},\bm{h})&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{(a)}\}}\displaystyle\int\mathcal{D}(z_{\mu,a})\exp\Bigg{[}\dfrac{\beta N}{2}\displaystyle\sum\limits_{\mu=1}^{L}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}\\ \\ &&+\sqrt{\beta N}\displaystyle\sum\limits_{\mu>L}^{K}\displaystyle\sum\limits_{a=1}^{L}m_{\mu}^{a}z_{\mu,a}\Bigg{]}\end{array}$ (A.4)
where $\mathcal{D}(z_{\mu,a})$ is the Gaussian measure with covariance $\bm{g}$. We compute the self-averaging statistical pressure $\mathcal{A}(\beta,H,\bm{g},\bm{h})$, defined as
@@
with the quenched expectation taken over the patterns $\xi^{\mu}$, by using the Guerra’s interpolation method, namely:
[…]
with $\frac{\mathcal{A}(\beta,H,\bm{g},\bm{h};t)}{dt}=\lim\limits_{N\to\infty}\frac{1}{N}\mathbb{E}\omega_{t}(\frac{\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{h};t)}{dt})\equiv\lim\limits_{N\to\infty}\frac{1}{N}\langle\frac{\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{h};t))}{dt}\rangle_{t}$, where we defined the quenched expectation over the (interpolating) Boltzmann average $\omega_{t}$ as
@@
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{h};t)&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{(a)}\}}\displaystyle\int\mathcal{D}(z_{\mu,a})\exp\Bigg{[}t\beta N\displaystyle\sum\limits_{\mu=1}^{L}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}+\sqrt{t}\sqrt{\dfrac{\beta}{N}}\displaystyle\sum\limits_{\mu>L,i=1}^{K,N}\displaystyle\sum\limits_{a=1}^{L}\xi_{i}^{\mu}\sigma_{i}^{a}z_{\mu,a}\\ \\ &&+(1-t)N\displaystyle\sum\limits_{\mu,a}\psi^{(a)}m_{\mu}^{a}+\sqrt{1-t}\displaystyle\sum\limits_{\mu}\tilde{Y}_{\mu}\displaystyle\sum\limits_{a}B^{(a)}z_{\mu,a}+\sqrt{1-t}\displaystyle\sum\limits_{i}Y_{i}\displaystyle\sum\limits_{a}A^{(a)}\sigma_{i}^{a}\\ \\ &&+\dfrac{1-t}{2}\displaystyle\sum\limits_{i}\displaystyle\sum\limits_{a\neq b}C^{(ab)}\sigma_{i}^{a}\sigma_{i}^{b}+\dfrac{1-t}{2}\displaystyle\sum\limits_{\mu}\displaystyle\sum\limits_{a\neq b}\tilde{C}^{(ab)}z_{\mu,a}z_{\mu,b}+\dfrac{1-t}{2}\displaystyle\sum\limits_{\mu}\displaystyle\sum\limits_{a}d^{a}(z_{\mu,a})^{2}\Bigg{]}\end{array}$ (A.8)
Setting the order parameters
$\begin{array}[]{lll}\bar{m}_{\mu}^{a}=\mathbb{E}\dfrac{1}{N}\displaystyle\sum\limits_{i}\xi_{i}^{\mu}\omega(\sigma_{i}^{a})&&\mathrm{with}\;\;\;a,\mu=1,\ldots,L\\ \\ q_{11}^{ab}=\dfrac{1}{N}\displaystyle\sum\limits_{i}\omega(\sigma_{i}^{a}\sigma_{i}^{b})&&q_{11}^{a}=1\\ \\ q_{12}^{ab}=\dfrac{1}{N}\displaystyle\sum\limits_{i}\omega(\sigma_{i}^{a})\omega(\sigma_{i}^{b})&&q_{12}^{a}=\dfrac{1}{N}\displaystyle\sum\limits_{i}\omega^{2}(\sigma_{i}^{a})\\ \\ p_{11}^{ab}=\dfrac{1}{K-L}\displaystyle\sum\limits_{\mu>L}\omega(z_{\mu,a}z_{\mu,b})&&p_{11}^{a}=\dfrac{1}{K-L}\displaystyle\sum\limits_{\mu>L}\omega((z_{\mu,a})^{2})\\ \\ p_{12}^{ab}=\dfrac{1}{K-L}\displaystyle\sum\limits_{\mu>L}\omega(z_{\mu,a})\omega(z_{\mu,b})&&p_{12}^{a}=\dfrac{1}{K-L}\displaystyle\sum\limits_{\mu>L}\omega^{2}(z_{\mu,a})\end{array}$ (A.9)
@@
the $t-$ derivative of $\mathcal{A}(\beta,H,\bm{g},\bm{h};t)$, after we have set the interpolating constants as
$\begin{array}[]{lll}(A^{(a)})^{2}=\beta\gamma\bar{p}_{12}^{a};&&A^{(a)}A^{(b)}=\beta\gamma\bar{p}_{12}^{ab}\\ \\ (B^{(a)})^{2}=\beta\bar{q}_{12}^{a};&&B^{(a)}B^{(b)}=\beta\bar{q}_{12}^{ab}\\ \\ C^{(a)}=\beta(1-\bar{q}_{12}^{a});&&\tilde{C}^{(ab)}=\beta(\bar{q}_{11}^{ab}-\bar{q}_{12}^{ab})\\ \\ C^{(ab)}=\beta\gamma(\bar{p}_{11}^{ab}-\bar{p}_{12}^{ab}).&&\end{array}$ (A.10)
where $\gamma=\lim\limits_{N\to\infty}K/N$, and under the RS assumption, reads
@@
$\frac{d\mathcal{A}(\beta,H,\bm{g},\bm{h};t)}{dt}=-\dfrac{\beta}{2}\displaystyle\sum\limits_{a,b=1}^{L}\displaystyle\sum\limits_{\mu=1}^{L}\bar{m}_{\mu}^{a}g_{ab}\bar{m}_{\mu}^{b}-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a\neq b}\Big{(}\bar{p}_{11}^{ab}\bar{q}_{11}^{ab}-\bar{p}_{12}^{ab}\bar{q}_{12}^{ab}\Big{)}-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a}\bar{p}_{12}^{a}\Big{(}1-\bar{q}_{12}^{a}\Big{)}.$ (A.11)
@@
Now we only need to compute the one-body term ($\mathcal{A}(\beta,H,\bm{g},\bm{h};t=0)$). We start form (A.8) setting $t=0$
@@
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{h};t=0)&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{(a)}\}}\displaystyle\int\mathcal{D}(z_{\mu,a})\exp\Bigg{[}\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}+\\ \\ &&+N\displaystyle\sum\limits_{\mu,a}\psi^{(a)}m_{\mu}^{a}+\displaystyle\sum\limits_{\mu}\tilde{Y}_{\mu}\displaystyle\sum\limits_{a}B^{(a)}z_{\mu,a}+\displaystyle\sum\limits_{i}Y_{i}\displaystyle\sum\limits_{a}A^{(a)}\sigma_{i}^{a}\\ \\ &&+\dfrac{1}{2}\displaystyle\sum\limits_{i}\displaystyle\sum\limits_{a\neq b}C^{(ab)}\sigma_{i}^{a}\sigma_{i}^{b}+\dfrac{1}{2}\displaystyle\sum\limits_{\mu}\displaystyle\sum\limits_{a\neq b}\tilde{C}^{(ab)}z_{\mu,a}z_{\mu,b}+\dfrac{1}{2}\displaystyle\sum\limits_{\mu}\displaystyle\sum\limits_{a}d^{a}(z_{\mu,a})^{2}\Bigg{]}\end{array}$ (A.12)
@@
$\mathcal{A}(\beta,H,\bm{g},\bm{h};t=0)=\lim_{N\to\infty}\frac{1}{N}\mathbb{E}\ln\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{\xi},\bm{h};t=0).$ (A.13)
@@
$\begin{array}[]{lll}\mathcal{A}(\beta,H,\bm{g},\bm{h};t=0)&=&\mathbb{E}_{\bm{\xi},x}\log\Bigg{\{}\displaystyle\sum\limits_{\{\bm{\sigma}^{(a)}\}}\exp\Bigg{(}\displaystyle\sum\limits_{a=1}^{L}\left[\displaystyle\sum\limits_{\mu=1}^{L}\beta\left(\bar{m}_{\mu}^{a}-\lambda\displaystyle\sum\limits_{b\neq a}\bar{m}_{\mu}^{b}\right)\xi^{\mu}+\beta Hh^{a}(t)+x\sqrt{\beta\gamma\bar{p}_{12}^{a}}\right]\sigma^{(a)}\\ \\ &&+\displaystyle\sum\limits_{a\neq b}\beta\gamma(\bar{p}_{11}^{ab}-\bar{p}_{12}^{ab})\sigma^{(a)}\sigma^{(b)}\Bigg{)}\Bigg{\}}\\ \\ &&-\dfrac{\gamma}{2}\log\Big{[}\mathrm{det}\,\mathcal{G}\Big{]}+\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\end{array}$ (A.14)

### v2

@@ 4.4 Checking disentanglement properties by Monte Carlo simulations
Figure 7: We consider a dataset consisting of 10 digits, each represented by $56\times 56$ pixels, see the upper part of the figure. This dataset is Hebbian-stored in a three-layer network governed by the cost function (5) with parameters $N=3136$, $K=10$, $H=0.1$, and $\lambda=0.19$. A mixture of digits $0,1,2$ is then prepared and presented as input to each layer of the network, which is subsequently updated through MC simulations with $\beta=1.9$. The lower part of the figure illustrates the evolution of the neuronal configurations $\bm{\sigma}^{1}$, $\bm{\sigma}^{2}$, and $\bm{\sigma}^{3}$. Correspondingly, the middle part displays the evolution of the associated Mattis magnetizations $\bm{m}^{1}$, $\bm{m}^{2}$, and $\bm{m}^{3}$. Different colors are used to distinguish the magnetizations related to the different patterns comprising the dataset, with emphasis on the digits included in the input mixture, as highlighted in the legend.
@@ 5 Conclusions
In particular, here, given a dataset of binary vectors $\bm{\xi}=\{\bm{\xi}^{\mu}\}_{\mu=1,...,K}\in\{-1,+1\}^{N\times K}$, the input is given by mixtures like $\bm{\sigma}^{(h)}=\textrm{sign}(\bm{\xi}^{1}+\bm{\xi}^{2}+...+\bm{\xi}^{L})$ – and this is interpreted as the initial neuronal configuration for each layer, that is, $\bm{\sigma}^{a}=\bm{\sigma}^{(h)}$ for a=1,,La=1,...,L – while the desired output is given by $\bm{\sigma}^{(1,2,...,L)}:\bm{\sigma}^{\ell}=\bm{\xi}^{\ell}$ for $\ell=1,...,L$ (without loss of generality) – and this is interpreted as the target state reached by the system.
@@
We emphasize that the kind of interactions implemented in this network yields a plethora of minima which can impair the disentanglement of the neuronal configuration $\bm{\sigma}^{(h)}$ into $\bm{\sigma}^{(1,2,...,L)}$. A way to see this is by considering an equivalent model obtained by applying a Hubbard-Stratonovich transformation to the model’s partition function (see App. A) and notice that the interaction among the dummy variables $z$’s is characterized by a high degree of frustration, especially compared with other layered associative-memory models, see e.g., AABCR-2024 . Many possible adjustments can be implemented to improve the performance of this model, for instance one can revise the Hebbian kernel to obtain a projection kernel Kohonen-1972 ; FAB-NN2019 that reduces the detrimental effects due to interference among the stored patterns, or allow for higher-order interactions FAB-NN2019 ; Krotov2018 ; Pedreschi1 ; Pedreschi2 which make the desired minima more stable. In fact, the architecture studied here opens several avenues. For instance, it would be natural to investigate the learning capabilities of Restricted Boltzmann Machines, which are equivalent to these modular networks (as highlighted in AABCR-2024 ), or possibly to extend the framework to spiking neural networks (see, e.g., Shen_Xu_Liu_Wang_Pan_Tang_2023 ; Shen_Ni_Xu_Tang_2024 ; CinesiIngordi ; xu2023enhancing ).
@@ Appendix A RS solution by interpolation technique
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{\xi},\bm{h})&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{1}\}}\cdots\displaystyle\sum\limits_{\{\bm{\sigma}^{L}\}}\exp\Bigg[\dfrac{\beta N}{2}\displaystyle\sum\limits_{\mu=1}^{K}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}\Bigg].\end{array}$ (A.2)
For completeness, we also recall the full list of observables:
$\begin{array}[]{lll}\bar{m}_{\mu}^{a}=\mathbb{E}\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\xi_{i}^{\mu}\omega(\sigma_{i}^{a})&&\mathrm{with}\;\;\;a,\mu=1,\ldots,L\\ \\ q_{11}^{ab}=\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\omega(\sigma_{i}^{a}\sigma_{i}^{b},)&&q_{11}^{a}=1,\\ \\ q_{12}^{ab}=\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\omega(\sigma_{i}^{a})\omega(\sigma_{i}^{b}),&&q_{12}^{a}=\dfrac{1}{N}\displaystyle\sum\limits_{i=1}^{N}\omega^{2}(\sigma_{i}^{a}),\\ \\ p_{11}^{ab}=\dfrac{1}{K-L}\displaystyle\sum\limits_{\mu>L}^{K}\omega(z_{\mu}^{a}z_{\mu}^{b}),&&p_{11}^{a}=\dfrac{1}{K-L}\displaystyle\sum\limits_{\mu>L}^{K}\omega((z_{\mu}^{a})^{2}),\\ \\ p_{12}^{ab}=\dfrac{1}{K-L}\displaystyle\sum\limits_{\mu>L}^{K}\omega(z_{\mu}^{a})\omega(z_{\mu}^{b})&&p_{12}^{a}=\dfrac{1}{K-L}\displaystyle\sum\limits_{\mu>L}^{K}\omega^{2}(z_{\mu}^{a}).\end{array}$ (A.3)
@@
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{\xi},\bm{h})&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{1}\}}\cdots\displaystyle\sum\limits_{\{\bm{\sigma}^{L}\}}\exp\Bigg[\dfrac{\beta N}{2}\displaystyle\sum\limits_{\mu=1}^{L}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\dfrac{\beta N}{2}\displaystyle\sum\limits_{\mu>L}^{K}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}\Bigg]\end{array}$ (A.4)
@@
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{\xi},\bm{h})&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{1}\}}\cdots\displaystyle\sum\limits_{\{\bm{\sigma}^{L}\}}\displaystyle\int\mathcal{D}(z)\exp\Bigg[\dfrac{\beta N}{2}\displaystyle\sum\limits_{\mu=1}^{L}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}\\ \\ &&+\sqrt{\beta N}\displaystyle\sum\limits_{\mu>L}^{K}\displaystyle\sum\limits_{a=1}^{L}m_{\mu}^{a}z_{\mu}^{a}\Bigg]\end{array}$ (A.5)
where $\mathcal{D}(z)$ is the Gaussian measure with covariance $\bm{g}^{-1}$. We compute the self-averaging statistical pressure $\mathcal{A}(\beta,H,\bm{g},\bm{h})$, defined as
@@
with the quenched expectation taken over the patterns $\xi^{\mu}$, by using the Guerra’s interpolation method. The basic idea to compute the free energy within this framework, is to introduce an interpolating parameter $t\in[0,1]$ and a corresponding interpolating free energy function $\mathcal{A}(\beta,H,\bm{g},\bm{h};t)$. This function is constructed so that at $t=1$, it coincides with the free energy of the original model, i.e., $\mathcal{A}(\beta,H,\bm{g},\bm{h};t=1)=\mathcal{A}(\beta,H,\bm{g},\bm{h})$. At the other endpoint, when $t=0$, $\mathcal{A}(\beta,H,\bm{g},\bm{h};t=0)$ corresponds to the free energy of a simplified one-body system, where each neuron is decoupled from the rest and instead interacts with a properly designed external field. This field is tailored to reproduce, at least in terms of low-order statistics, the internal field that would be generated by the actual network.
In defining the one-body system (i.e., at $t=0$), we also include auxiliary constants and functions, which we retain the flexibility to choose later in a way that simplifies the analysis and, in the thermodynamic limit ($N\to\infty$), ensures Replica Symmetry (RS). Under the RS ansatz, we assume that the probability distributions of the order parameters become Dirac deltas in the thermodynamic limit, hence the expectations of the order parameters collapse on these values in this asymptotic limit, that is, calling $x$ a generic order parameter, $\lim_{N\to\infty}\langle x(\bm{\sigma})\rangle=\bar{x}$.
Ultimately, the central mathematical tool used here is the Fundamental Theorem of Calculus, which serves as a natural link between the two boundary cases of the interpolating parameter. This leads us to the sum rule that follows:
$\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h})=\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};t=0)+\int_{0}^{1}dt\>\frac{\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};s)}{ds}\Bigg|_{s=t},$ (A.7)
with $\frac{\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};t)}{dt}=\lim\limits_{N\to\infty}\frac{1}{N}\mathbb{E}\omega_{t}(\frac{\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{h};t)}{dt})\equiv\lim\limits_{N\to\infty}\frac{1}{N}\langle\frac{\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{h};t))}{dt}\rangle_{t}$, where we defined the quenched expectation over the (interpolating) Boltzmann average $\omega_{t}$ as
@@
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{h};t)&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{1}\}}\cdots\displaystyle\sum\limits_{\{\bm{\sigma}^{L}\}}\displaystyle\int\mathcal{D}(z)\exp\Bigg[t\beta N\displaystyle\sum\limits_{\mu=1}^{L}\displaystyle\sum\limits_{a,b=1}^{L}m_{\mu}^{a}g_{ab}m_{\mu}^{b}+\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}+\sqrt{t}\sqrt{\dfrac{\beta}{N}}\displaystyle\sum\limits_{\mu>L,i=1}^{K,N}\displaystyle\sum\limits_{a=1}^{L}\xi_{i}^{\mu}\sigma_{i}^{a}z_{\mu}^{a}\\ \\ &&+(1-t)N\displaystyle\sum\limits_{a=1}^{L}\displaystyle\sum\limits_{\mu=1}^{L}\psi^{(a)}m_{\mu}^{a}+\sqrt{1-t}\displaystyle\sum\limits_{\mu>L}^{K}\tilde{Y}_{\mu}\displaystyle\sum\limits_{a=1}^{L}B^{(a)}z_{\mu}^{a}+\sqrt{1-t}\displaystyle\sum\limits_{i=1}^{N}Y_{i}\displaystyle\sum\limits_{a=1}^{L}A^{(a)}\sigma_{i}^{a}\\ \\ &&+\dfrac{1-t}{2}\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{\begin{subarray}{c}a,b=1\\ a\neq b\end{subarray}}^{L}C^{(ab)}\sigma_{i}^{a}\sigma_{i}^{b}+\dfrac{1-t}{2}\displaystyle\sum\limits_{\mu>L}^{K}\displaystyle\sum\limits_{\begin{subarray}{c}a,b=1\end{subarray}}^{L}\tilde{C}^{(ab)}z_{\mu}^{a}z_{\mu}^{b}\Bigg].\end{array}$ (A.9)
@@
the $t$- derivative of $\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};t)$, after we have set the interpolating constants as
[…]
where $\gamma=\lim\limits_{N\to\infty}K/N$, can be written as
@@
$\frac{d\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};t)}{dt}=-\dfrac{\beta}{2}\displaystyle\sum\limits_{a,b=1}^{L}\displaystyle\sum\limits_{\mu=1}^{L}\bar{m}_{\mu}^{a}g_{ab}\bar{m}_{\mu}^{b}-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{\begin{subarray}{c}a,b=1\\ a\neq b\end{subarray}}^{L}\Big(\bar{p}_{11}^{ab}\bar{q}_{11}^{ab}-\bar{p}_{12}^{ab}\bar{q}_{12}^{ab}\Big)-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a=1}^{L}\bar{p}_{12}^{a}\Big(1-\bar{q}_{12}^{a}\Big).$ (A.11)
@@
Now we only need to compute the one-body term ($\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};t=0)$). We start form (A.9) setting $t=0$
@@
$\begin{array}[]{lll}\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{h};t=0)&=&\displaystyle\sum\limits_{\{\bm{\sigma}^{1}\}}\cdots\displaystyle\sum\limits_{\{\bm{\sigma}^{L}\}}\displaystyle\int\mathcal{D}(z)\exp\Bigg[\beta H\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{a=1}^{L}h_{i}^{a}(t)\sigma_{i}^{a}+\\ \\ &&+N\displaystyle\sum\limits_{\mu=1}^{L}\displaystyle\sum\limits_{a=1}^{L}\psi^{(a)}m_{\mu}^{a}+\displaystyle\sum\limits_{\mu>L}^{K}\tilde{Y}_{\mu}\displaystyle\sum\limits_{a=1}^{L}B^{(a)}z_{\mu}^{a}+\displaystyle\sum\limits_{i=1}^{N}Y_{i}\displaystyle\sum\limits_{a=1}^{L}A^{(a)}\sigma_{i}^{a}\\ \\ &&+\dfrac{1}{2}\displaystyle\sum\limits_{i=1}^{N}\displaystyle\sum\limits_{\begin{subarray}{c}a,b=1\\ a\neq b\end{subarray}}^{L}C^{(ab)}\sigma_{i}^{a}\sigma_{i}^{b}+\dfrac{1}{2}\displaystyle\sum\limits_{\mu>L}^{K}\displaystyle\sum\limits_{a,b=1}^{L}\tilde{C}^{(ab)}z_{\mu}^{a}z_{\mu}^{b}\Bigg]\end{array}$ (A.12)
@@
$\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};t=0)=\lim_{N\to\infty}\frac{1}{N}\mathbb{E}\ln\mathcal{Z}_{N}(\beta,H,\bm{g},\bm{\xi},\bm{h};t=0).$ (A.13)
@@
$\begin{array}[]{lll}\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};t=0)&=&\mathbb{E}_{\bm{\xi},x}\log\Bigg\{\displaystyle\sum\limits_{\{\bm{\sigma}^{a}\}}\exp\Bigg(\displaystyle\sum\limits_{a=1}^{L}\left[\displaystyle\sum\limits_{\mu=1}^{L}\beta\left(\bar{m}_{\mu}^{a}-\lambda\displaystyle\sum\limits_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{L}\bar{m}_{\mu}^{b}\right)\xi^{\mu}+\beta Hh^{a}(t)+x\sqrt{\beta\gamma\bar{p}_{12}^{a}}\right]\sigma^{a}\\ \\ &&+\displaystyle\sum\limits_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{L}\beta\gamma(\bar{p}_{11}^{ab}-\bar{p}_{12}^{ab})\sigma^{a}\sigma^{(b)}\Bigg)\Bigg\}\\ \\ &&-\dfrac{\gamma}{2}\log\Big[\mathrm{det}\,\mathcal{G}\Big]+\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\end{array}$ (A.14)